Assets 2017 2018 2018 Current assets Cash Accounts receivable Inventory Total $ 24,046 12,448 25,392 $ 61,886 $ 24, 255 15,23Prepare the 2018 statement of cash flows for Earl Grey Golf Corp. (Negative answers should be indicated by a minus sign. OmitNet cash from investment activities Financing activities Net cash from financing activities Cash, end of year
Answer #1

Answer:

CASH FLOW STATEMENT:

A B C D
1 Particulars Amount ($) Amount ($)
2 Cash Beginning of the Year 24,046
3 Operating Activities:
4 Net Income 43,780
5 ADD Depreciation 32,220
6 ADD Increase in Accounts Payable 4,236 [=27,300-23,384]
7 ADD Increase in other Liabilities 3,982 [=16,700-11,771]
8 LESS Increase in Accounts Receivable 2,787 [=12,448-15,235]
9 LESS Increase in Inventory 1,763 [=25,392-27,155]
10
11 Net Cash From Operating Activities 79,668 [=SUM(D2:D10)]
12 Investing Activities:
13 LESS Fixed Assets Acquisition 73,259 [=-365,734+324,695-32,220]
14
15 Nat Cash From Investing Activities -73,259 [=D13]
16 Financing Activities:
17 Decrease in Notes Payable -1,200 [=10,800-12,000]
18 Dividend Paid -20,000
19 Increase in long term debt 15,000 [95,000-80,000]
20
21 Net Cash From Financing Activities -6,200 [=SUM(D17:D19)]
22
23 Net Increase in Cash 209 [=D11+D15+D21]
24 Cash,End of the Year 24,255 [=D2+D23]